Calculate the t value using SPSS. Please show your SPSS output. You should have 3 tables:...

  1. Calculate the t value using SPSS. Please show your SPSS output. You should have 3 tables: Paired Samples Statistics, Paired Samples Correlations, and Paired Samples Test.

Pretest

Posttest

Difference

D2

3

4

-1

1

2

7

-5

25

4

6

-2

4

2

4

-2

4

1

5

-4

16

5

8

-3

9

6

9

-3

9

3

8

-5

25

2

6

-4

16

3

7

-4

16

Sum = 31

Sum = 64

Sum = -33

Sum = 125

Mean = 3.1

Mean = 6.4

Mean = -3.3

Mean – 12.5
Answer #1

Answer:-

Given that:-

Paired Samples Statistics, Paired Samples Correlations, and Paired Samples Test..

Given Sample size n = 10\For the score difference we have

\overline{D} = -3.3

S_{D} = 1.337

Null and Alternative Hypothesis:

H_{0} : \mu_{D} = 0

H_{a} : \mu_{D} \neq 0

H_{0} : The two means are Same

H_{a} : The two means are different

Test Statistic:

t = \frac{\overline{D}}{\frac{S_{D}}{\sqrt{n}}}

=  \frac{-3.3}{\frac{1.337}{\sqrt{10}}}

= -7.802

\alpha = 0.05, df = 9

t_{c} = 2.262 (using t table)

Since that \vert t\vert = 7.802 > t_{c} = 2.262, we Reject H_{0}

\therefore There is enough euiclence to claim that population mean  \mu_{1} is different than  \mu_{2}.